# More Agents Is All You Need

### v1

@@ Abstract
We find that, simply via a sampling-and-voting method, the performance of large language models (LLMs) scales with the number of agents instantiated. Also, this method is orthogonal to existing complicated methods to further enhance LLMs, while the degree of enhancement is correlated to the task difficulty. We conduct comprehensive experiments on a wide range of LLM benchmarks to verify the presence of our finding, and to study the properties that can facilitate its occurrence. Our code is publicly available at: Git.
Machine Learning, ICML
@@ 1 Introduction
In these works, multiple LLM agents are used to improve the performance of LLMs. For instance, LLM-Debate (Du et al., 2023) employs multiple LLM agents in a debate form. The reasoning performance is improved by creating a framework that allows more than one agent to “debate” the final answer of arithmetic tasks. They show performance improvements compared to using one single agent. Similarly, CoT-SC (Wang et al., 2023b) generates multiple thought chains and picks the most self-consistent one as the final answer. The reasoning performance is improved by involving more thought chains compared to chain-of-thought (CoT) (Wei et al., 2022) which employs a single thought chain. Incidentally, from the data analysis of these works, we can notice the effects of putting multiple agents together, to some extent, can lead to a performance improvement in certain problems. For example, in Table 10 of Section 3.3 of LLM-Debate (Du et al., 2023), the authors have reported a preliminary curve: the accuracy of a math problem increases with the number of debating agents (although the number was simply increased from 1 to 7). Also, in Wang et al. (2023b), involving more chains-of-thought pipelines (termed as a “sample-and-marginalize” decoding procedure), can lead to a performance gain. We realize that the LLM performance may likely be improved by a brute-force scaling up the number of agents instantiated. However, since the scaling property of “raw” agents is not the focus of these works, the scenarios/tasks and experiments considered are limited. So far, there lacks a dedicated in-depth study on this phenomenon. Hence, a natural question arises: Does this phenomenon generally exist?
Figure 1: The accuracy increases with ensemble size across Llama2-13B, Llama2-70B and GPT-3.5-Turbo in GSM8K. When the ensemble size scales up to $15$, Llama2-13B achieves comparable accuracy with Llama2-70B. Similarly, When the ensemble size scales up to $15$ and $20$, Llama2-70B and GPT-3.5-Turbo achieve comparable accuracy with their more powerful counterparts.
To answer the research question above, we conduct the first comprehensive study on the scaling property of LLM agents. To dig out the potential of multiple agents, we propose to use a simple(st) sampling-and-voting method, which involves two phases. First, the query of the task, i.e., the input to an LLM, is iteratively fed into a single LLM, or a multiple LLM-Agents collaboration framework, to generate multiple outputs. Subsequently, majority voting is used to determine the final result. The procedure is inspired by that of the CoT-SC, but it does not rely on designing complex CoT paths. In fact, it can be used as a plug-in to further enhance CoT-based methods, as will be shown in our evaluations.
@@
Additionally, the experimental results indicate a correlation between the efficacy of the performance improvements and the difficulty of the problems addressed. To understand the reasons behind these performance improvements, we analyze the influence of problem difficulty on the effectiveness of our method. We classify difficulty into three dimensions: the inherent difficulty, the length of reasoning steps, and the prior probability of the correct answer. Through a series of experiments, we adjust these dimensions and observe their effects independently. We observe and summarize a few properties, based on which, we further develop optimization strategies that can intrigue the power of “More Agents”.
@@ 2 Related Work
Related works can be categorized into three parts: 1) LLM self-ensemble (Wang et al., 2023b), which attempts to harness multiple outputs from homogeneous LLMs to assemble the final answer; 2) heterogeneous LLM ensemble, which focuses on combining heterogeneous LLMs through supervised learning to improve performance across various downstream applications; and 3) multiple LLM agents collaboration, which improves performance through interactions among LLM agents. We discuss these works below.
LLM Self-Ensemble. CoT-SC (Wang et al., 2023b) harnesses diverse chain-of-thought (Wei et al., 2022) prompts to elicit a variety of reasoning processes from a single LLM and select the final answer through majority voting. Fu et al. (2023); Li et al. (2023b); Cobbe et al. (2021b); Thoppilan et al. (2022); Lin et al. (2023) can be considered as the extensions of CoT-SC. These methods mainly focus on reasoning tasks and exclusively investigate the compatibility with CoT. In contrast, our method not only validates effectiveness in reasoning tasks but also in generation tasks. Moreover, our method is compatible with a broader range of methods, such as prompt engineering (including CoT) and multiple LLM agents collaboration. Very recently, Lu et al. (2024) proposes a method named Blended that utilizes multiple LLMs for chat scenarios. In contrast, Blended focuses on utilizing the power of multiple LLMs, whereas our focus is on the scaling trend of adding more LLMs. Also, Blended is only for limited chat scenarios evaluated via human annotations. Furthermore, we explore orthogonality with other methods.
@@
Multiple LLM Agents Collaboration. Studies explore various multiple LLM agents interaction architectures, with employing static debate-style engagements among LLMs for enhanced reasoning (Du et al., 2023; Liang et al., 2023; Xiong et al., 2023). Liu et al. (2023) enables agents to interact for multiple rounds in a dynamic architecture. Li et al. (2023a); Hong et al. (2023); Wu et al. (2023); Chen et al. (2023c, a) offer several multi-agent frameworks that enable the development of LLM applications or enhance task-solving capabilities. However, these methods primarily focus on the interaction structures between LLM agents, rather than the relationship between the number of agents and performance. We also select representative methods (Du et al., 2023; Shinn et al., 2023) to combine with our method, achieving further enhancements.
@@ 3 Method
Figure 2: Illustration of our method. The two-phase process begins by feeding the task query, either alone or combined with prompt engineering methods, into LLM agents to generate answers. Subsequently, majority voting is applied to these answers to determine the final answer. Specifically, an LLM agent refers to a single LLM or a multiple LLM-Agents collaboration framework.
In this section, we introduce our method which is implemented through a two-phase process: sampling and voting. The overview of our method is shown in Figure 2.
@@
Voting. Let $A$ represent the final answer. In this phase, we employ majority voting to consolidate the response sample set $S$ into the final answer $A$. This involves calculating the cumulative similarity for each sample relative to the others, denoted as $V(s_{i})=\sum_{j=1,j\neq i}^{N}sim(s_{i},s_{j})$. For open-ended generation tasks such as code generation, the BLEU score (Papineni et al., 2002) is utilized to quantify similarity. Conversely, for close-ended tasks like multiple-choice questions, similarity is measured by occurrence frequency. The sample that exhibits the highest cumulative similarity is then chosen as the final answer denoted as $A=\operatorname*{arg\,max}_{s_{i}\in S}V(s_{i})$.
The complete process of the sampling-and-voting method is described in Algorithm 15.
Algorithm 1 Sampling-and-voting
@@ 4 Experimental Setup
Methods Various LLMs Tasks Integrated with Methods Chat Arithmetic Reasoning General Reasoning Code Generation Prompt Engineering Multiple LLM-Agents Collaboration CoT-SC (Wang et al., 2023b) Only CoT (Wei et al., 2022) Complexity-CoT (Fu et al., 2023) Only CoT (Wei et al., 2022) Debate (Du et al., 2023) Blended (Lu et al., 2024) Ours
@@
  • Arithmetic Reasoning. Similar to Wang et al. (2023b); Fu et al. (2023); Du et al. (2023), we select the GSM8K (Cobbe et al., 2021a) as one of the test sets. Additionally, we select the more challenging MATH dataset (Hendrycks et al., 2021b), which is used by Wu et al. (2023).
  • General Reasoning. Similar to Du et al. (2023); Jiang et al. (2023), we select the MMLU (Hendrycks et al., 2021a). Additionally, we select the dataset from the chess state tracking task (Chess)  [^1] Chess State Tracking, which is used by Du et al. (2023); Zhang et al. (2023).
  • Code Generation. Similar to Liu et al. (2023), we select the HumanEval (Chen et al., 2021). To implement our method, we compute the BLEU score (Papineni et al., 2002) among all pairs of generated candidate answers. The answer with the highest cumulative BLEU score is then selected as the final output.
@@
We evaluate our method using language models of different scales from the Llama2 (Touvron et al., 2023) and GPT series (OpenAI, 2022). Specifically, we evaluate two versions of Llama2-Chat [^2] Llama2-Chat, optimized for conversational use cases through alignment techniques, with model sizes of 13B and 70B parameters. Additionally, we include GPT-3.5-Turbo and GPT-4 in our evaluation.
@@
  • Prompt Engineering. Various prompt engineering methods are considered to conduct comprehensive experiments. We evaluate Chain-of-Thought prompting (CoT) (Wei et al., 2022), Zero-Shot Chain-of-Thought prompting (Zero-Shot Cot) (Kojima et al., 2022), and more sophisticated methods such as Solo Performance Prompting (SPP) (Wang et al., 2023c). Initially, these methods are applied with a single LLM query. We then increase the number of queries and employ majority voting to determine the most consistent answer as the final response.
  • Multiple LLM Agents Collaboration. We select LLM-Debate (Du et al., 2023) denoted as Debate, and self-reflection (Shinn et al., 2023) denoted as Reflection. Within these methods, we generate multiple samples by iteratively operating these methods and using majority voting to produce the final answer.
Specifically, the effectiveness of our method is evaluated by averaging the results across $10$ independent runs. During each run, we scale up the ensemble size to $40$ to ensure maximum gains. However, when integrating our method with the Debate (Du et al., 2023), the ensemble size is limited to $10$ due to the significant computational overhead introduced by the communication architecture. Detailed experimental settings are provided in the Appendix A.
@@ 5 Experimental Results
Figure 3: The accuracy scales with the ensemble size of our method across different tasks with various LLMs.
Table 2: Our method generally enhances performance across all tasks and LLMs. The bolded instances indicate that smaller LLMs outperform the larger LLMs. “Single” denotes that the LLM is queried only once. GPT-4 is used only for comparison with other methods, hence it only presents “Single” results.
Model GSM8K MATH Chess MMLU HumanEval Single Ours Single Ours Single Ours Single Ours Single Ours Llama2-13B (Touvron et al., 2023) 0.35 0.59 (+0.24) 0.03 0.09 (+0.06) 0.14 0.18 (+0.04) 0.42 0.51 (+0.09) 0.14 0.18 (+0.04) Llama2-70B (Touvron et al., 2023) 0.54 0.74 (+0.20) 0.05 0.11 (+0.06) 0.12 0.13 (+0.01) 0.55 0.60 (+0.05) 0.24 0.33 (+0.09) GPT-3.5-Turbo (OpenAI, 2022) 0.73 0.85 (+0.12) 0.29 0.39 (+0.10) 0.51 0.55 (+0.04) 0.59 0.70 (+0.11) 0.67 0.73 (+0.06) GPT-4 (OpenAI, 2022) 0.88 0.40 0.65 0.77 0.88
@@
Model Method GSM8K MATH Chess MMLU HumanEval Standalone +Ours Standalone +Ours Standalone +Ours Standalone +Ours Standalone +Ours Llama2-13B (Touvron et al., 2023) COT (Wei et al., 2022) 0.39 0.56 (+0.17) 0.04 0.06 (+0.02) 0.18 0.23 (+0.07) 0.42 0.43 (+0.01) 0.13 0.20 (+0.07) ZS-COT (Kojima et al., 2022) 0.40 0.61 (+0.21) 0.03 0.08 (+0.05) 0.15 0.20 (+0.05) 0.42 0.48 (+0.06) 0.15 0.22 (+0.07) SPP (Wang et al., 2023c) 0.19 0.42 (+0.23) 0.01 0.04 (+0.03) 0.21 0.26 (+0.05) 0.32 0.53 (+0.21) 0.03 0.08 (+0.05) Debate (Du et al., 2023) 0.38 0.48 (+0.10) 0.05 0.07 (+0.02) 0.18 0.19 (+0.01) 0.37 0.39 (+0.02) 0 0 Reflection (Shinn et al., 2023) 0.36 0.59 (+0.23) 0.01 0.03 (+0.02) 0.13 0.19 (+0.06) 0.45 0.50 (+0.05) 0.06 0.13 (+0.07) Ours 0.59 0.09 0.18 0.51 0.25 Llama2-70B (Touvron et al., 2023) COT (Wei et al., 2022) 0.57 0.72 (+0.15) 0.06 0.13 (+0.07) 0.10 0.11 (+0.01) 0.56 0.57 (+0.01) 0.30 0.32 (+0.02) ZS-COT (Kojima et al., 2022) 0.57 0.73 (+0.16) 0.04 0.10 (+0.06) 0.20 0.27 (+0.07) 0.54 0.65 (+0.11) 0.23 0.29 (+0.06) SPP (Wang et al., 2023c) 0.42 0.69 (+0.27) 0.03 0.09 (+0.06) 0.16 0.27 (+0.11) 0.49 0.63 (+0.14) 0.15 0.20 (+0.05) Debate (Du et al., 2023) 0.59 0.65 (+0.06) 0.10 0.11 (+0.01) 0.14 0.17 (+0.03) 0.56 0.58 (+0.02) 0 0 Reflection (Shinn et al., 2023) 0.52 0.77 (+0.25) 0.02 0.05 (+0.03) 0.15 0.26 (+0.11) 0.42 0.55 (+0.13) 0.16 0.26 (+0.10) Ours 0.74 0.11 0.13 0.60 0.33 GPT-3.5-Turbo (OpenAI, 2022) COT (Wei et al., 2022) 0.74 0.84 (+0.10) 0.28 0.41 (+0.13) 0.50 0.55 (+0.05) 0.61 0.64 (+0.03) 0.70 0.75 (+0.05) ZS-COT (Kojima et al., 2022) 0.74 0.88 (+0.14) 0.25 0.40 (+0.15) 0.35 0.48 (+0.13) 0.58 0.69 (+0.11) 0.67 0.74 (+0.07) SPP (Wang et al., 2023c) 0.70 0.83 (+0.13) 0.26 0.39 (+0.13) 0.37 0.54 (+0.17) 0.53 0.68 (+0.15) 0.57 0.64 (+0.07) Debate (Du et al., 2023) 0.83 0.85 (+0.02) 0.32 0.36 (+0.04) 0.49 0.57 (+0.08) 0.56 0.67 (+0.11) 0.18 0.24 (+0.06) Reflection (Shinn et al., 2023) 0.76 0.84 (+0.08) 0.27 0.41 (+0.14) 0.44 0.57 (+0.13) 0.39 0.44 (+0.05) 0.58 0.73 (+0.15) Ours 0.85 0.39 0.55 0.70 0.73
Method +Ours GPT-3.5 70B 13B Overall COT (Wei et al., 2022) 2.8 3.6 3.6 3.3 ZS-COT (Kojima et al., 2022) 2.8 2.4 3 2.7 SPP (Wang et al., 2023c) 4.6 3.6 3.8 4 Debate (Du et al., 2023) 3.8 4.4 5 4.4 Reflection (Shinn et al., 2023) 3 4.0 3 3.3 Ours 2.6 2.6 2.2 2.5
Figure 4: Our method improves accuracy over various hyperparameters and tasks.

### v2

# algorithmic

@@ Abstract
We find that, simply via a sampling-and-voting method, the performance of large language models (LLMs) scales with the number of agents instantiated. Also, this method, termed as Agent Forest, is orthogonal to existing complicated methods to further enhance LLMs, while the degree of enhancement is correlated to the task difficulty. We conduct comprehensive experiments on a wide range of LLM benchmarks to verify the presence of our finding, and to study the properties that can facilitate its occurrence. Our code is publicly available at: https://github.com/MoreAgentsIsAllYouNeed/AgentForest.
@@ 1 Introduction
In these works, multiple LLM agents are used to improve the performance of LLMs. For instance, LLM-Debate Du et al. (2023) employs multiple LLM agents in a debate form. The reasoning performance is improved by creating a framework that allows more than one agent to “debate” the final answer of arithmetic tasks. They show performance improvements compared to using one single agent. Similarly, CoT-SC (Wang et al., 2023b) generates multiple thought chains and picks the most self-consistent one as the final answer. The reasoning performance is improved by involving more thought chains compared to chain-of-thought (CoT) (Wei et al., 2022) which employs a single thought chain. Incidentally, from the data analysis of these works, we can notice the effects of putting multiple agents together, to some extent, can lead to a performance improvement in certain problems. For example, in Table 10 of Section 3.3 of LLM-Debate Du et al. (2023), the authors have reported a preliminary curve: the accuracy of a math problem increases with the number of debating agents (although the number was simply increased from 1 to 7). Also, in Wang et al. (2023b), involving more chain-of-thought pipelines (termed as a “sample-and-marginalize” decoding procedure), can lead to a performance gain. We realize that the LLM performance may likely be improved by a brute-force scaling up of the number of agents instantiated. However, since the scaling property of “raw” agents is not the focus of these works, the scenarios/tasks and experiments considered are limited. So far, there lacks a dedicated in-depth study on such a phenomenon. Hence, a natural question arises: Does this phenomenon generally exist?
Figure 1: The accuracy increases with ensemble size across Llama2-13B, Llama2-70B and GPT-3.5-Turbo in GSM8K. When the ensemble size scales up to $15$, Llama2-13B achieves comparable accuracy with Llama2-70B. Similarly, When the ensemble size scales up to $15$ and $20$, Llama2-70B and GPT-3.5-Turbo achieve comparable accuracy with their more powerful counterparts. The error bars represent the standard error.
To answer the research question above, we conduct the first comprehensive study on the scaling property of LLM agents. To dig out the potential of multiple agents, we propose to use a simple(st) sampling-and-voting method, which involves two phases. First, the query of the task, i.e., the input to an LLM, is iteratively fed into a single LLM, or a multiple LLM-Agents collaboration framework, to generate multiple outputs. Subsequently, majority voting is used to determine the final result. The procedure is inspired by that of the CoT-SC, but it does not rely on designing complex CoT paths. In fact, it can be used as a plug-in to further enhance CoT-based methods, as will be shown in our evaluations. Our method is termed as Agent Forest, a tribute to the classic Random Forest (Breiman, 2001).
@@
Additionally, the experimental results indicate that there are greater performance improvements when addressing difficult tasks and when using weaker models. To understand the reasons behind these performance improvements, we analyze the influence of problem difficulty on the effectiveness of our method. We classify difficulty into three dimensions: the inherent difficulty, the length of reasoning steps, and the prior probability of the correct answer. Through a series of experiments, we adjust these dimensions and observe their effects independently. We observe and summarize a few properties, based on which, we further develop optimization strategies that can intrigue the power of “More Agents”.
@@ 2 Related Work
Related works can be categorized into three parts: 1) LLM self-ensemble Wang et al. (2023b), which attempts to harness multiple outputs from homogeneous LLMs to assemble the final answer; 2) heterogeneous LLM ensemble, which focuses on combining heterogeneous LLMs through supervised learning to improve performance across various downstream applications; and 3) multiple LLM agents collaboration, which improves performance through interactions among LLM agents. We discuss these works below.
LLM Self-Ensemble. CoT-SC Wang et al. (2023b) harnesses diverse chain-of-thought Wei et al. (2022) prompts to elicit a variety of reasoning processes from a single LLM and select the final answer through majority voting. Fu et al. (2023); Li et al. (2023b); Cobbe et al. (2021b); Thoppilan et al. (2022); Lin et al. (2023) can be considered as the extensions of CoT-SC. These methods mainly focus on reasoning tasks and exclusively investigate the compatibility with CoT. In contrast, our method not only validates effectiveness in reasoning tasks but also in generation tasks. Moreover, our method is compatible with a broader range of methods, such as prompt engineering (including CoT) and multiple LLM agents collaboration. Very recently, Lu et al. (2024) proposes a method named Blended that utilizes multiple LLMs for chat scenarios. In contrast, Blended focuses on utilizing the power of multiple LLMs, whereas our focus is on the scaling trend of adding more LLMs. Also, Blended is only for limited chat scenarios evaluated via human annotations. Furthermore, we explore orthogonality with other methods.
@@
Multiple LLM Agents Collaboration. Du et al. (2023); Liang et al. (2023); Xiong et al. (2023) explore various multiple LLM agents interaction architectures, with employing static debate-style engagements among LLMs for enhanced reasoning . Liu et al. (2023) enables agents to interact for multiple rounds in a dynamic architecture. Li et al. (2023a); Hong et al. (2023); Wu et al. (2023); Chen et al. (2023c; a) offer several multi-agent frameworks that enable the development of LLM applications or enhance task-solving capabilities. However, these methods primarily focus on the interaction structures between LLM agents, rather than the relationship between the number of agents and performance. We also select representative methods Du et al. (2023); Shinn et al. (2023) to combine with our method, achieving further enhancements.
@@ 3 Method
Figure 2: Illustration of Agent Forest. The two-phase process begins by feeding the task query, either alone or combined with prompt engineering methods, into LLM agents to generate answers. Subsequently, majority voting is applied to these answers to determine the final answer. Specifically, an LLM agent refers to a single LLM or a multiple LLM-Agents collaboration framework.
In this section, we introduce Agent Forest, which is implemented through a two-phase process: sampling and voting. The overview of our method is shown in Figure 2.
@@
Voting. Let $A$ represent the final answer. In this phase, we employ majority voting to consolidate the response sample set $S$ into the final answer $A$. This involves calculating the cumulative similarity for each sample relative to the others, denoted as $V(s_{i})=\sum_{j=1,j\neq i}^{N}sim(s_{i},s_{j})$. For open-ended generation tasks such as code generation, the BLEU score proposed by Papineni et al. (2002) is utilized to quantify similarity. Conversely, for close-ended tasks like multiple-choice questions, similarity is measured by occurrence frequency. The sample that exhibits the highest cumulative similarity is then chosen as the final answer denoted as $A=\operatorname*{arg\,max}_{s_{i}\in S}V(s_{i})$.
The complete process of Agent Forest is described in Algorithm 15.
Algorithm 1 Agent Forest
@@ 4 Experimental Setup
  • Arithmetic Reasoning. Similar to Wang et al. (2023b); Fu et al. (2023); Du et al. (2023), we select the GSM8K Cobbe et al. (2021a) as one of the test sets. Additionally, we select the more challenging MATH dataset Hendrycks et al. (2021b), which is used by Wu et al. (2023).
  • General Reasoning. Similar to Du et al. (2023); Jiang et al. (2023), we select the MMLU Hendrycks et al. (2021a). Additionally, we select the dataset from the chess state tracking task (Chess)  [^1] Chess State Tracking, which is used by Du et al. (2023); Zhang et al. (2023).
  • Code Generation. Similar to Liu et al. (2023), we select the HumanEval Chen et al. (2021). To implement our method, we compute the BLEU score Papineni et al. (2002) among all pairs of generated candidate answers. The answer with the highest cumulative BLEU score is then selected as the final output.
Methods Various LLMs Tasks Integrated with Methods Chat Arithmetic Reasoning General Reasoning Code Generation Prompt Engineering Multiple LLM-Agents Collaboration CoT-SC Wang et al. (2023b) Only CoT Wei et al. (2022) Complexity-CoT Fu et al. (2023) Only CoT Wei et al. (2022) Debate Du et al. (2023) Blended Lu et al. (2024) Ours
@@
We evaluate our method using language models of different scales from the Llama2 Touvron et al. (2023) and GPT series OpenAI (2022). Specifically, we evaluate two versions of Llama2-Chat [^2] Llama2-Chat, optimized for conversational use cases through alignment techniques, with model sizes of 13B and 70B parameters. Additionally, we include GPT-3.5-Turbo and GPT-4 in our evaluation.
@@
  • Prompt Engineering. Various prompt engineering methods are considered to conduct comprehensive experiments. We evaluate Chain-of-Thought prompting (CoT) Wei et al. (2022), Zero-Shot Chain-of-Thought prompting (Zero-Shot Cot) Kojima et al. (2022), and more sophisticated methods such as Solo Performance Prompting (SPP) Wang et al. (2023c). Initially, these methods are applied with a single LLM query. We then increase the number of queries and employ majority voting to determine the most consistent answer as the final response.
  • Multiple LLM Agents Collaboration. We select LLM-Debate Du et al. (2023) denoted as Debate, and self-reflection Shinn et al. (2023) denoted as Reflection. Within these methods, we generate multiple samples by iteratively operating these methods and using majority voting to produce the final answer.
Specifically, the effectiveness of our method is evaluated by averaging the results across $10$ independent runs. During each run, we scale up the ensemble size to $40$ to ensure maximum gains. However, when integrating our method with the Debate Du et al. (2023), the ensemble size is limited to $10$ due to the significant computational overhead introduced by the communication architecture. Detailed experimental settings are provided in the Appendix A.
@@ 5 Experimental Results
Figure 3: The accuracy scales with the ensemble size of our method across different tasks with various LLMs. The error bars represent the standard error.
Table 2: Our method generally enhances performance across all tasks and LLMs. The bolded instances indicate that smaller LLMs outperform the larger LLMs. “Single” denotes that the LLM is queried only once. GPT-4 is used only for comparison with other methods, hence it only presents “Single” results. “Ours” denotes our method where the ensemble size is 40. The error bars represent the standard error.
Model GSM8K MATH Chess MMLU HumanEval Single Ours Single Ours Single Ours Single Ours Single Ours Llama2-13B Touvron et al. (2023) 0.35 ± 3e-2 0.59 ± 5e-4 0.03 ± 7e-3 0.09 ± 2e-3 0.14 ± 2e-2 0.18 ± 2e-3 0.42 ± 3e-2 0.51 ± 1e-3 0.14 ± 1e-2 0.18 ± 1e-3 Llama2-70B Touvron et al. (2023) 0.54 ± 3e-2 0.74 ± 1e-3 0.05 ± 1e-2 0.11 ± 1e-3 0.12 ± 2e-2 0.13 ± 2e-3 0.55 ± 2e-2 0.60 ± 3e-3 0.24 ± 1e-2 0.33 ± 1e-3 GPT-3.5-Turbo OpenAI (2022) 0.73 ± 2e-2 0.85 ± 3e-3 0.29 ± 2e-2 0.39 ± 2e-3 0.51 ± 3e-2 0.55 ± 2e-3 0.59 ± 3e-2 0.70 ± 2e-3 0.67 ± 2e-2 0.73 ± 1e-2 GPT-4 OpenAI (2022) 0.88 ± 2e-2 0.40 ± 3e-2 0.65 ± 2e-2 0.77 ± 2e-2 0.88 ± 3e-2
@@
Model Method GSM8K MATH Chess MMLU HumanEval Standalone +Ours Standalone +Ours Standalone +Ours Standalone +Ours Standalone +Ours Llama2-13B Touvron et al. (2023) COT Wei et al. (2022) 0.39 0.56 (+0.17) 0.04 0.06 (+0.02) 0.18 0.23 (+0.07) 0.42 0.43 (+0.01) 0.13 0.20 (+0.07) ZS-COT Kojima et al. (2022) 0.40 0.61 (+0.21) 0.03 0.08 (+0.05) 0.15 0.20 (+0.05) 0.42 0.48 (+0.06) 0.15 0.22 (+0.07) SPP Wang et al. (2023c) 0.19 0.42 (+0.23) 0.01 0.04 (+0.03) 0.21 0.26 (+0.05) 0.32 0.53 (+0.21) 0.03 0.08 (+0.05) Debate Du et al. (2023) 0.38 0.48 (+0.10) 0.05 0.07 (+0.02) 0.18 0.19 (+0.01) 0.37 0.39 (+0.02) 0 0 Reflection Shinn et al. (2023) 0.36 0.59 (+0.23) 0.01 0.03 (+0.02) 0.13 0.19 (+0.06) 0.45 0.50 (+0.05) 0.06 0.13 (+0.07) Ours 0.59 0.09 0.18 0.51 0.25 Llama2-70B Touvron et al. (2023) COT Wei et al. (2022) 0.57 0.72 (+0.15) 0.06 0.13 (+0.07) 0.10 0.11 (+0.01) 0.56 0.57 (+0.01) 0.30 0.32 (+0.02) ZS-COT Kojima et al. (2022) 0.57 0.73 (+0.16) 0.04 0.10 (+0.06) 0.20 0.27 (+0.07) 0.54 0.65 (+0.11) 0.23 0.29 (+0.06) SPP Wang et al. (2023c) 0.42 0.69 (+0.27) 0.03 0.09 (+0.06) 0.16 0.27 (+0.11) 0.49 0.63 (+0.14) 0.15 0.20 (+0.05) Debate Du et al. (2023) 0.59 0.65 (+0.06) 0.10 0.11 (+0.01) 0.14 0.17 (+0.03) 0.56 0.58 (+0.02) 0 0 Reflection Shinn et al. (2023) 0.52 0.77 (+0.25) 0.02 0.05 (+0.03) 0.15 0.26 (+0.11) 0.42 0.55 (+0.13) 0.16 0.26 (+0.10) Ours 0.74 0.11 0.13 0.60 0.33 GPT-3.5-Turbo OpenAI (2022) COT Wei et al. (2022) 0.74 0.84 (+0.10) 0.28 0.41 (+0.13) 0.50 0.55 (+0.05) 0.61 0.64 (+0.03) 0.70 0.75 (+0.05) ZS-COT Kojima et al. (2022) 0.74 0.88 (+0.14) 0.25 0.40 (+0.15) 0.35 0.48 (+0.13) 0.58 0.69 (+0.11) 0.67 0.74 (+0.07) SPP Wang et al. (2023c) 0.70 0.83 (+0.13) 0.26 0.39 (+0.13) 0.37 0.54 (+0.17) 0.53 0.68 (+0.15) 0.57 0.64 (+0.07) Debate Du et al. (2023) 0.83 0.85 (+0.02) 0.32 0.36 (+0.04) 0.49 0.57 (+0.08) 0.56 0.67 (+0.11) 0.18 0.24 (+0.06) Reflection Shinn et al. (2023) 0.76 0.84 (+0.08) 0.27 0.41 (+0.14) 0.44 0.57 (+0.13) 0.39 0.44 (+0.05) 0.58 0.73 (+0.15) Ours 0.85 0.39 0.55 0.70 0.73
Figure 4: Our method improves accuracy over various hyperparameters and tasks. The default $T$ is 1.0 and the default $p$ is 1.0.